OpenStudy (anonymous):

Find the derivative of f(x) = 4 divided by x at x = 2.

OpenStudy (sidsiddhartha):

use \[\frac{ d }{ dx }[x^n]=nx^{n-1}\]

OpenStudy (anonymous):

I haven't learned that yet (Pre-calc) We use the Difference Quotient

OpenStudy (sidsiddhartha):

all right so take \[f(x)=4/x\] so \[f(x+h)=4/(x+h)\] ok?

OpenStudy (anonymous):

ok

OpenStudy (sidsiddhartha):

\[f'(x)=\lim_{h \rightarrow 0}\frac{ f(x+h)-f(x) }{ h}\] familiar with this?

OpenStudy (anonymous):

yes, we subtract 4/x, divide all by h

OpenStudy (sidsiddhartha):

ok now just put values of f(x) and f(x+h) inside the limit \[f'(x)=\lim_{h \rightarrow 0}\frac{ 4 }{ h }[\frac{ 1 }{ (x+h) }-\frac{ 1 }{ x }]\] ok??

OpenStudy (anonymous):

ok

OpenStudy (sidsiddhartha):

now just simplify it \[f'(x)=\lim_{h \rightarrow 0}\frac{ 4 }{ h }[\frac{ x-x-h }{ (x+h)x}]\] \[f'(x)=\lim_{h \rightarrow 0}[\frac{ -4 }{ x(x+h)}]=\frac{ -4 }{ x^2}\]

OpenStudy (sidsiddhartha):

got this?

OpenStudy (anonymous):

yes

OpenStudy (sidsiddhartha):

now just put x=2 so it will be\[f'(2)=\frac{ -4 }{ 2^2}=-1\]

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

Thanks

OpenStudy (sidsiddhartha):

np :)